OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

Do you know how to find the slope of the line containing both of those points?

OpenStudy (anonymous):

i think y1-y2/x1-x2

OpenStudy (anonymous):

That's it, so what is the slope you get from that?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Let's look at the y's first: y2 = 3 and y1 = 2 y2-y1= 1 Now the x's x2 = 6 x1 = 2 x2-x1= 4 So our slope will be: m = 1/4

OpenStudy (anonymous):

From here we use the point slope formula. We can use either point they'll give you the same equation at the end. The point slope formula is: y−y1=m(x−x1)

OpenStudy (anonymous):

So: m = 1/4 x1 = 6 y1 = 2 Putting that into our formula: y - 2 = 1/4 (x - 6) Now we solve for y: y - 2 = 1/4 x - 6/4 y = 1/4 x + 2 - 6/4 y = 1/4 x + 2/4 y = 1/4 x + 1/2

OpenStudy (anonymous):

4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6 these are the answers i got the first wrong and i do not knw where i went wrong